# Sprecher Networks: A Parameter-Efficient Kolmogorov-Arnold Architecture

### v1

@@ 8.2 Impact of lateral mixing
To avoid confounds from implementation-level vectorization (e.g., a highly vectorized KAN forward vs. a less vectorized SN implementation), we do not treat wall-clock time as a primary figure of merit here. We record and report it for completeness, but treat it as secondary and highly implementation-dependent; we therefore do not draw algorithmic conclusions from wall-clock time.
Structure of this subsection. We present each comparison task as its own subsubsection, with (i) a brief problem description, (ii) a one-paragraph explanation for why the task might reward SN inductive biases (if applicable), and (iii) a compact result table for that task.
8.3.1 Monotone single-index with quantiles (MQSI; 20D, m=9m{=}9 heads)
Setup.
@@ 8.3.1 Monotone single-index with quantiles (MQSI; 20D, m=9m{=}9 heads)
We sample inputs $\mathbf{x}\sim\mathrm{Unif}([0,1]^{D})$ with $D=20$ and define $m=9$ heads at quantile levels $\tau_{j}=\mathrm{linspace}(0.1,0.9)$, with $z_{j}=\Phi_{\mathcal{N}}^{-1}(\tau_{j})=\sqrt{2}\,\mathrm{erf}^{-1}(2\tau_{j}-1)$, where $\Phi_{\mathcal{N}}$ is the standard normal CDF (to avoid confusion with the outer spline $\Phi$). For each coordinate, define a monotone sigmoid feature
@@ Discussion.
$y_{j}(x)\;=\;\tanh\!\Big(\beta\big[\mu(x)+\sigma(x)z_{j}+a_{\mathrm{shift}}\,q_{j}+q_{\mathrm{bias}}\,q_{j}\big]\Big),$
with $x\in[0,1]^{D}$, $j=0,\dots,m{-}1$, $q_{j}\in[-1,1]$ on a uniform grid, and $z_{j}=\Phi_{\mathcal{N}}^{-1}(\tau_{j})$. Here we set $D=12$ and $m=64$ and construct $\mu(\mathbf{x})$ and $\sigma(\mathbf{x})$ as follows. For each coordinate,
@@
Setup. D=12D{=}12, m=64m{=}64, $a_{\mathrm{shift}}{=}0.7$, $\beta{=}0.8$, $q_{\mathrm{bias}}{=}0.6$; train for 4000 epochs on CPU on a fixed i.i.d. training set of $n_{\mathrm{train}}{=}1024$ samples; $n_{\mathrm{test}}{=}50{,}000$. BN evaluation uses per-batch statistics while freezing the running-statistics buffers (i.e., no running-mean/variance updates during evaluation). Parameter parity was enforced by choosing the KAN basis size $K$ to be as close as possible to (and not exceeding) the SN parameter count, yielding #params $\approx 3072$ for both models; the resulting KAN basis size was K=4K{=}4. The specific model configurations were:
  • SN: Two hidden Sprecher blocks of width $32$ (architecture $12\to 32\to 32\to 64$), with $60$ knots for each inner spline $\phi$ and each outer spline $\Phi$; BatchNorm after each block except the first; linear residual connections; no lateral mixing; and dynamic spline-domain updates for the first $400$ epochs, frozen thereafter.
@@
Summary across baseline tasks.
Table 6 collates the head-to-heads above. In both cases we observe SNs outperform KANs at matched parameter budgets and training epochs. The MQSI task, which directly matches the SN compositional/shift bias, shows the clearest gap in correlation-structure fidelity; the DenseHeadShift task is closer, with SNs typically ahead by a modest margin.
@@
Table 6: Summary across baseline tasks at matched parameter counts and $4000$ epochs; wall-clock time shown for completeness. Times are per-task averages across seeds.
@@
  • Epochs/optimizer: $4000$ epochs for both models. We use Adam with gradient-norm clipping $1.0$. For SNs we use learning rate $3\times 10^{-4}$ for the main parameters, $5\times 10^{-4}$ for lateral-mixing parameters (when enabled), and $10^{-3}$ for $(c_{c},c_{r})$ codomain parameters (when enabled); weight decay $10^{-7}$. For KANs we use learning rate $10^{-3}$ and weight decay $10^{-6}$ (no other tuning beyond the fairness knobs).
  • Residuals/BN: Linear residuals in both; BatchNorm layers placed “after” each layer/block, and evaluated using per-batch statistics without updating running-mean/variance buffers.
  • Parity: Parameter matching via a closed-loop count of KAN parameters as a function of $K$ (prefer $\leq$).
  • Data: Identical $(x_{\text{train}},y_{\text{train}})$ and test sets across models per task; fixed seeds.
Takeaways.
These controlled baselines demonstrate that there exist natural tasks, particularly those exhibiting shared-index families or smooth saturations, on which SNs achieve lower test error than parameter-matched KANs under equal training budgets. The observed margins range from modest to substantial (especially on structured multi-head problems); we expect this picture to sharpen as we expand the suite with $\sim$8 additional tasks in the same format.

### v2

@@ 8.2 Impact of lateral mixing
We report test RMSE, parameter counts, and (where available) wall-clock training time. We do not report train RMSE and we do not treat wall-clock time as a primary metric (it is sensitive to implementation details and hardware).
Structure of this subsection. We present eight barebones comparisons (§8.3.1) across diverse regression tasks designed to isolate architectural differences between SNs and KANs.
8.3.1 Barebones comparisons (no residuals, no normalization)
In these experiments, we strip away all engineering enhancements and focus on the fundamental architectural differences between SNs and KANs:
  • SN: Piecewise-linear (PWL) splines; dynamic spline-domain updates during the first 10% of training (400/4000 epochs), then frozen; no residuals; no normalization; no lateral mixing; no learned $\Phi$-codomain parameters.
  • KAN: Cubic Hermite (PCHIP) splines with fixed uniform knot grids; no grid updates; no residuals; no normalization; linear extrapolation outside the knot interval.
The target functions are chosen to highlight settings where the SN’s structural inductive biases, particularly shared inner splines across input dimensions and systematic head shifts, provide an architectural advantage. For most benchmarks, these advantages stem from the Sprecher structure itself (parameter sharing, shifted-input reuse); we also include one benchmark (PWL-vs-PCHIP) that tests spline-type flexibility on non-smooth targets. Table 8 summarizes the results; we describe each benchmark below.
Benchmark Params (SN/KAN) SN RMSE KAN RMSE Ratio SN Time (s) KAN Time (s)
Softstair-Wavepacket (10D) 2761/2761 $\mathbf{0.352\pm 0.003}$ $0.551\pm 0.009$ $1.57\times$ $38.4\pm 5.9$ $56.3\pm 7.8$
InputShift-Bump (12D, 17h) 1599/1597 $\mathbf{0.253\pm 0.001}$ $0.339\pm 0.105$ $1.34\times$
Shared-Warped-Ridge (16D) 2115/2113 $\mathbf{0.701\pm 0.002}$ $0.805\pm 0.100$ $1.15\times$ $266.7\pm 64.2$ $250.2\pm 97.6$
Shared-Warp-Chirp (10D) 2447/2449 $\mathbf{0.896\pm 0.025}$ $1.470\pm 0.060$ $1.64\times$ $549.5\pm 107.0$ $432.7\pm 124.9$
Motif-Chirp (10D) 2106/2107 $\mathbf{0.123\pm 0.006}$ $0.182\pm 0.015$ $1.48\times$ $305.3\pm 100.4$ $259.0\pm 89.5$
Oscillatory-HeadShift (12D, 64h) 2763/2768 $\mathbf{0.669\pm 0.020}$ $0.752\pm 0.016$ $1.12\times$ $366.3\pm 39.0$ $225.1\pm 28.3$
PWL-vs-PCHIP (10D) 1701/1701 $\mathbf{0.732\pm 0.012}$ $0.998\pm 0.035$ $1.36\times$ $76.5\pm 11.4$ $71.6\pm 6.4$
Quantile Harmonics (2D) 2026/2025 $\mathbf{0.395\pm 0.146}$ $0.619\pm 0.035$ $1.57\times$ $1236.1\pm 82.4$ $3790.8\pm 489.3$
Table 8: Barebones SN vs. KAN comparisons (20 seeds each; no residuals, no normalization, no lateral mixing; and no learned $\Phi$-codomain parameters in SN). RMSE values are mean $\pm$ std across seeds; bold indicates the winner. Ratio == KAN/SN (higher means larger SN advantage). Unless otherwise noted, RMSE is the final-epoch test RMSE; Benchmark 3 reports the best test RMSE observed during training, and Benchmark 8 reports test RMSE at the best validation checkpoint. Wall-clock times (in seconds) are mean $\pm$ std where per-model timings are available. SN uses PWL splines with dynamic spline-domain updates during the first 10% of training (400/4000 epochs), then frozen; KAN uses cubic PCHIP splines with fixed grids.
Benchmark 1: Softstair-Wavepacket (10D).
The first benchmark tests a smooth but highly non-unimodal, permutation-symmetric target built from a shared monotone coordinate warp and a localized oscillatory wavepacket. For $t\in[0,1]$, define
$\mathrm{softstair}(t)=t+0.25\,\operatorname{sigmoid}(25(t-0.20))+0.35\,\operatorname{sigmoid}(35(t-0.55))+0.20\,\operatorname{sigmoid}(60(t-0.85)).$
For $\mathbf{x}\in[0,1]^{10}$ let $s(\mathbf{x})=\frac{1}{10}\sum_{i=1}^{10}\mathrm{softstair}(x_{i})$ and
$y(\mathbf{x})=0.70\,\exp\!\bigl(-30(s(\mathbf{x})-0.90)^{2}\bigr)\Bigl[\sin(14\pi s(\mathbf{x}))+0.30\sin(42\pi s(\mathbf{x}))\Bigr]+0.05\,(s(\mathbf{x})-0.90).$
With $n_{\text{train}}=2048$ and $n_{\text{test}}=8192$ i.i.d. uniform samples, SN achieves $0.352\pm 0.003$ RMSE while KAN obtains $0.551\pm 0.009$ (ratio $1.57\times$). The target applies the same 1D warp to each coordinate before pooling, matching SN’s shared inner spline; a KAN must allocate separate edge splines per coordinate to learn this shared warp under the same parameter budget.
Benchmark 2: InputShift-Bump (12D, 17 heads).
This multi-output regression defines $y(\mathbf{x})\in\mathbb{R}^{17}$ with components indexed by $q\in\{0,\dots,16\}$ through a head-dependent input shift. Let
$\phi_{\star}(t)=\operatorname{sigmoid}(8(t-0.5)),\qquad\Phi_{\star}(s)=\exp\!\left(-\frac{1}{2}\left(\frac{s-8.0}{2.6}\right)^{2}\right)+0.25\sin(0.85\,s),$
fix a mixing vector $w\in\mathbb{R}^{12}$ (sampled once per seed and then held fixed). Concretely, sample $\tilde{w}\sim\mathcal{N}(0,I_{12})$, set $\tilde{w}\leftarrow\tilde{w}/\|\tilde{w}\|$, let $s\in\mathbb{R}^{12}$ have entries $s_{i}=(-1)^{i-1}$, and set $w=0.9\,\tilde{w}+0.1\,s$, and define
$s_{q}(\mathbf{x})=q+\sum_{i=1}^{12}w_{i}\,\phi_{\star}(x_{i}+0.045\,q),\qquad y_{q}(\mathbf{x})=\Phi_{\star}(s_{q}(\mathbf{x}))+\varepsilon_{q},$
with i.i.d. Gaussian noise $\varepsilon_{q}\sim\mathcal{N}(0,0.25^{2})$ per sample/head. Trained on $n_{\text{train}}=1024$ and tested on $n_{\text{test}}=50000$, SN obtains $0.253\pm 0.001$ RMSE while KAN obtains $0.339\pm 0.105$ ($1.34\times$). The very large KAN variance suggests sensitivity to the head-conditioned shift inside $\phi_{\star}$; SN’s explicit $x_{i}\mapsto x_{i}+\eta q$ structure and spline-domain updates during an initial phase (first $400$ of $4000$ epochs, then frozen) appear to improve stability across heads.
Benchmark 3: Shared-Warped-Ridge (16D).
Here we evaluate a symmetric shared-warp ridge. Define
$h(t)=0.6\,t+0.2\,\operatorname{sigmoid}(30(t-0.30))+0.2\,\operatorname{sigmoid}(30(t-0.70)),\qquad s(\mathbf{x})=\frac{1}{16}\sum_{i=1}^{16}h(x_{i}),$
and
$y(\mathbf{x})=\sin(12\pi s^{2})\,\exp(-3(s-0.5)^{2})+0.15\sin(2\pi s)+0.10(s-0.5).$
With $n_{\text{train}}=1024$, $n_{\text{test}}=8192$ and comparable parameters (2115 vs. 2113), SN achieves $0.701\pm 0.002$ (best test RMSE observed during training) while KAN’s $0.805\pm 0.100$ includes one outlier seed (seed 7 reaches $1.13$). This task matches SN’s inductive bias directly: the ground truth is exactly a shared 1D warp followed by a scalar post-map.
Benchmark 4: Shared-Warp-Chirp (10D).
This benchmark applies a shared monotone warp to each coordinate and then evaluates a chirp in the pooled coordinate. For $\mathbf{x}\in[0,1]^{10}$ define
$s(\mathbf{x})=\frac{1}{10}\sum_{i=1}^{10}\operatorname{sigmoid}(18(x_{i}-0.5)),\qquad y(\mathbf{x})=\sin\!\bigl(2\pi(6s(\mathbf{x})+5s(\mathbf{x})^{2})\bigr)+0.35\cos(2\pi\cdot 3s(\mathbf{x}))+0.10(s(\mathbf{x})-0.5).$
With $n_{\text{train}}=1024$, $n_{\text{test}}=8192$ and comparable parameters (2447 vs. 2449), SN achieves $0.896\pm 0.025$ while KAN obtains $1.470\pm 0.060$ ($1.64\times$). The quadratic phase term yields an increasing instantaneous frequency; under a fixed per-edge spline budget, SN benefits from learning the shared warp once.
Benchmark 5: Motif-Chirp (10D).
The motif benchmark is generated by a depth-2 Sprecher-style composition with internal widths $w_{1}=16$ and $w_{2}=15$. Let
$\phi_{\star}(t)=\operatorname{sigmoid}(30(t-0.5)),\qquad\Phi^{(1)}_{\star}(s)=\sin(1.7\,s)+0.28\,\bigl|\sin(0.55\,s+0.2)\bigr|,$
$\tilde{\phi}_{\star}(t)=\operatorname{sigmoid}(12t),\qquad\Phi^{(2)}_{\star}(s)=\sin(1.15\,s)+0.22\,\bigl|\sin(0.9\,s+0.7)\bigr|.$
Define deterministic positive weights by
$\tilde{\lambda}^{(1)}_{i}=0.30+0.70|\cos(0.8(i-1)+0.1)|,\quad i=1,\dots,10,\qquad\lambda^{(1)}_{i}=9\,\tilde{\lambda}^{(1)}_{i}/\sum_{j=1}^{10}\tilde{\lambda}^{(1)}_{j}$,
$\tilde{\lambda}^{(2)}_{q}=0.25+0.75|\sin(0.6q+0.4)|,\quad q=0,\dots,w_{1}-1,\qquad\lambda^{(2)}_{q}=7\,\tilde{\lambda}^{(2)}_{q}/\sum_{j=0}^{w_{1}-1}\tilde{\lambda}^{(2)}_{j}$.
For $q\in\{0,\dots,w_{1}-1\}$ set $s^{(1)}_{q}(\mathbf{x})=q+\sum_{i=1}^{10}\lambda^{(1)}_{i}\,\phi_{\star}(x_{i}+0.06q)$ and $h_{q}(\mathbf{x})=\Phi^{(1)}_{\star}(s^{(1)}_{q}(\mathbf{x}))$. For $r\in\{0,\dots,w_{2}-1\}$ set $s^{(2)}_{r}(\mathbf{x})=r+\sum_{q=0}^{w_{1}-1}\lambda^{(2)}_{q}\,\tilde{\phi}_{\star}(h_{q}(\mathbf{x})+0.05r)$, and finally
$y(\mathbf{x})=5\left(\frac{1}{w_{2}}\sum_{r=0}^{w_{2}-1}\Phi^{(2)}_{\star}(s^{(2)}_{r}(\mathbf{x}))\right)-0.4.$
With $n_{\text{train}}=2048$ and $n_{\text{test}}=8192$, SN achieves $0.123\pm 0.006$ while KAN reaches $0.182\pm 0.015$ ($1.48\times$). This task is literally a two-block Sprecher composition, so SN’s shared $(\phi,\Phi)$ pairs align closely with the ground-truth structure.
Benchmark 6: Oscillatory-HeadShift (12D, 64 heads).
This multi-head benchmark is a deterministic teacher $f:[0,1]^{12}\to\mathbb{R}^{64}$ with shared latent projections and a head-dependent phase shift. Train/test inputs are sampled i.i.d. uniformly over $0,1]^{12}$, and there is no additive observation noise. We sample the teacher parameters once (fixed seed 20240518) and hold them fixed across all runs: sample $\mathbf{w}_{1},\mathbf{w}_{2}\sim\mathcal{N}(0,I_{12})$ and normalize $\mathbf{w}_{j}\leftarrow\mathbf{w}_{j}/\|\mathbf{w}_{j}\|$ for $j\in\{1,2\}$, and sample $M\in\mathbb{R}^{12\times 3}$ with i.i.d. entries $\mathcal{N}(0,1/12)$. For $\mathbf{x}\in[0,1]^{12}$ let $\mathbf{x}_{c}=\mathbf{x}-\tfrac{1}{2}\mathbf{1}$, $u=\mathbf{w}_{1}^{\top}\mathbf{x}_{c}$, $v=\mathbf{w}_{2}^{\top}\mathbf{x}_{c}$, and $\mathbf{z}=\mathbf{x}_{c}M=(z_{1},z_{2},z_{3})\in\mathbb{R}^{3}$. For head index $h\in\{0,\dots,63\}$ define
$\tilde{h}=\frac{h-\tfrac{1}{2}(63)}{63},\qquad\delta_{h}=0.35\,\tilde{h},\qquad\omega=6\pi,$
$\theta^{(1)}_{h}=\omega(u+\delta_{h}),\qquad\theta^{(2)}_{h}=\tfrac{1}{2}\omega(v-0.7\,\delta_{h}),\qquad a(\mathbf{x})=0.75+0.25\tanh(2.5\,z_{1}),$
and output
$y_{h}(\mathbf{x})=a(\mathbf{x})\sin(\theta^{(1)}_{h})+0.35\cos(2\theta^{(2)}_{h})+0.15\sin(3\theta^{(1)}_{h}+0.25z_{2})+0.10\cos(\theta^{(2)}_{h}+0.35z_{3}).$
Trained on $n_{\text{train}}=1024$ and tested on $n_{\text{test}}=50000$, SN obtains $0.669\pm 0.020$ vs. KAN’s $0.752\pm 0.016$ ($1.12\times$). Here the gap is smaller but still consistent; SN exhibits slightly higher variance ($0.020$ vs. $0.016$), suggesting that this highly oscillatory setting is harder to optimize reliably for both models.
Benchmark 7: PWL-vs-PCHIP (10D).
This benchmark isolates a different (and arguably orthogonal) advantage: SN’s inner functions are piecewise-linear, whereas our KAN baseline uses C1 cubic PCHIP splines on edges. The target is a triangle wave applied to the coordinate mean, with a small affine term $y=\mathrm{tri}(12\cdot\bar{x})+0.05(\bar{x}-0.5)$, where $\bar{x}=\tfrac{1}{d}\sum_{i}x_{i}$ and $\mathrm{tri}(u)=2\left|2(u-\lfloor u\rfloor)-1\right|-1$ is a period-1 triangle wave in [1,1][-1,1], introducing cusps that a cubic spline must approximate. With $n_{\text{train}}=2048$ and $n_{\text{test}}=8192$, SN obtains $0.732\pm 0.012$ while KAN reaches $0.998\pm 0.035$ ($1.36\times$). The result suggests that even when the shared structure is trivial (a coordinate mean) and the dominant challenge is non-smoothness, the combination of (i) PWL splines and (ii) domain adaptation via spline-domain updates during an initial phase (first $400$ of $4000$ epochs, then frozen) can be a strong practical advantage.
Benchmark 8: Quantile Harmonics (2D).
This is a 2D scalar regression task with soft regime-switching along the $x$ coordinate (the “quantile” terminology refers to the soft bins, not to multi-quantile outputs). For $(x,y)\in[0,1]^{2}$, define four Gaussian gates
$g_{k}(x)=\exp\!\left(-\frac{1}{2}\left(\frac{x-c_{k}}{0.055}\right)^{2}\right),\quad c=(0.10,0.30,0.55,0.80),\quad\tilde{g}_{k}(x)=g_{k}(x)/\sum_{j}g_{j}(x),$
an amplitude modulation $a(x,y)=0.65+0.35\cos\!\bigl(2\pi(2y+0.15\sin(2\pi x))\bigr)$, and regime-specific harmonic mixtures
[…]
The target is
$t(x,y)=\sum_{k=1}^{4}\tilde{g}_{k}(x)\,m_{k}(x,y)+0.15\,\sin(2\pi(x+y))\exp(-6(y-0.5)^{2}).$
With $n_{\text{train}}=4096$ ($n_{\text{val}}=2048$) and $n_{\text{test}}=2048$, we report test RMSE at the best validation checkpoint: SN achieves $0.395\pm 0.146$ while KAN obtains $0.619\pm 0.035$ ($1.57\times$). The large SN variance indicates that optimization can occasionally fail; nevertheless SN wins on average and achieves substantially better median performance.
Barebones takeaways.
Across all eight benchmarks, the SN consistently outperforms the parameter-matched KAN, with win ratios ranging from $1.12\times$ to $1.64\times$. The SN’s variance across seeds is typically much lower than the KAN’s, indicating more reliable optimization (though Oscillatory-HeadShift and Quantile Harmonics are exceptions where SN shows higher variance). The strongest advantages appear on tasks whose ground truth reuses a shared 1D transform across many coordinates and/or across systematically shifted heads (Benchmarks 1–6), matching SNs’ shared-spline and head-shift parameterization. Benchmarks 5 and 7 also include mild non-smoothness ($|\sin|$ motifs and triangle-wave cusps), where the PWL-vs-cubic spline choice can contribute under a matched parameter budget. Benchmark 8 is a smooth 2D regime-switching harmonic mixture; SN remains better on average but shows occasional optimization failures reflected in higher variance.
8.3.2 Comparisons with linear residuals and BatchNorm
The following two benchmarks use linear residual connections and BatchNorm for both models, representing a more “production-like” configuration. While the barebones comparisons above isolate core architectural differences, these residual-equipped comparisons show that SN advantages persist even when both models benefit from standard deep learning enhancements.
MQSI Setup (20D, m=9m{=}9 heads).
@@ 8.3.2 Comparisons with linear residuals and BatchNorm
We sample inputs $\mathbf{x}\sim\mathrm{Unif}([0,1]^{D})$ with $D=20$ and define $m=9$ heads at quantile levels $\tau_{j}=\mathrm{linspace}(0.1,0.9)$, with $z_{j}=F_{\mathcal{N}}^{-1}(\tau_{j})=\sqrt{2}\,\mathrm{erf}^{-1}(2\tau_{j}-1)$, where $F_{\mathcal{N}}$ denotes the standard normal CDF. For each coordinate, define a monotone sigmoid feature
@@ Discussion.
$y_{j}(x)\;=\;\tanh\!\Big(\beta\big[\mu(x)+\sigma(x)z_{j}+a_{\mathrm{head}}\,q_{j}\big]\Big),\qquad a_{\mathrm{head}}:=a_{\mathrm{shift}}+q_{\mathrm{bias}},$
with $x\in[0,1]^{D}$, $j=0,\dots,m{-}1$, $q_{j}\in[-1,1]$ on a uniform grid (note: $q_{j}$ here denotes the task’s head-index coordinate, distinct from the SN output index $q$), and $z_{j}=\Phi_{\mathcal{N}}^{-1}(\tau_{j})$. Here we set $D=12$ and $m=64$ and construct $\mu(\mathbf{x})$ and $\sigma(\mathbf{x})$ as follows. For each coordinate,
@@
Setup. D=12D{=}12, m=64m{=}64, $a_{\mathrm{shift}}{=}0.7$, $\beta{=}0.8$, $q_{\mathrm{bias}}{=}0.6$ (so $a_{\mathrm{head}}{=}a_{\mathrm{shift}}+q_{\mathrm{bias}}{=}1.3$); train for 4000 epochs on CPU on a fixed i.i.d. training set of $n_{\mathrm{train}}{=}1024$ samples; $n_{\mathrm{test}}{=}50{,}000$. BN evaluation uses per-batch statistics while freezing the running-statistics buffers (i.e., no running-mean/variance updates during evaluation). Parameter parity was enforced by choosing the KAN basis size $K$ to be as close as possible to (and not exceeding) the SN parameter count, yielding #params $\approx 3072$ for both models; the resulting KAN basis size was K=4K{=}4. The specific model configurations were:
  • SN: Two hidden Sprecher blocks of width $32$ (architecture $12\to[32,32]\to 64$), with $60$ knots for each inner spline $\phi$ and each outer spline $\Phi$; BatchNorm after each block except the first; linear residual connections; no lateral mixing; and dynamic spline-domain updates during the first 10% of training (400/4000 epochs), then frozen.
@@
Residual-equipped summary.
Table 9 collates the residual-equipped comparisons. In both cases, SNs outperform KANs at matched parameter budgets and training epochs, even when both models benefit from linear residuals and BatchNorm. The MQSI task shows a clearer gap in correlation-structure fidelity; the DenseHeadShift task is closer, with SNs typically ahead by a modest margin.
@@
Table 9: Residual-equipped SN vs. KAN comparisons (linear residuals + BatchNorm for both; 10 seeds each). Times are per-task averages.
@@
  • Epochs/optimizer: $4000$ epochs for both models. Barebones (Table 8): Adam optimization with learning rate $10^{-3}$ for both models (Benchmark 3 uses $2\times 10^{-3}$ with cosine annealing); weight decay 0 except Benchmark 3 ($10^{-6}$); gradient-norm clipping is enabled in Benchmarks 2, 4, and 6 at threshold $1.0$ (both models) and in Benchmark 3 with max-norm $10.0$. Residual-equipped (Table 9): Adam with gradient-norm clipping $1.0$. SN learning rate $3\times 10^{-4}$; weight decay $10^{-7}$. KAN learning rate $10^{-3}$; weight decay $10^{-6}$.
  • Barebones: No residuals, no BatchNorm, no lateral mixing. SN uses PWL splines with dynamic spline-domain updates during the first 10% of training (400/4000 epochs), then frozen. KAN uses cubic PCHIP splines with fixed uniform knots.
  • Residual-equipped: Linear residuals in both; BatchNorm after each layer/block (except first), evaluated using per-batch statistics.
  • Parity: Parameter matching via closed-loop count of KAN parameters as a function of $K$ (prefer $\leq$).
  • Data: Identical $(x_{\text{train}},y_{\text{train}})$ and test sets across models per task; inputs are sampled from $\mathrm{Unif}([0,1]^{d})$, using i.i.d. draws except in low-dimensional visualization-oriented settings where we use deterministic uniform grids (1D: linspace; 2D: $\sqrt{n}\times\sqrt{n}$ mesh when $n$ is a perfect square). Dataset sizes: Benchmark 1 $n_{\text{train}}=2048$, $n_{\text{test}}=8192$; Benchmarks 2 and 6 $n_{\text{train}}=1024$, $n_{\text{test}}=50000$; Benchmarks 3 and 4 $n_{\text{train}}=1024$, $n_{\text{test}}=8192$; Benchmarks 5 and 7 $n_{\text{train}}=2048$, $n_{\text{test}}=8192$; Benchmark 8 $n_{\text{train}}=4096$, $n_{\text{val}}=2048$, $n_{\text{test}}=2048$. We use 20 seeds for barebones, 10 seeds for residual-equipped.
Overall takeaways.
Across both barebones and residual-equipped comparisons (ten benchmarks total, covering scalar and multi-head regression in 2D–20D), SNs consistently achieve lower test error than parameter-matched KANs under equal training budgets. The barebones comparisons (Table 8) demonstrate that the SN’s architectural advantages—shared inner splines, systematic head shifts, Sprecher-style composition, and piecewise-linear spline flexibility—provide genuine inductive biases across diverse task types. The residual-equipped comparisons (Table 9) show these advantages persist when both models use standard deep learning techniques. Win ratios range from $1.12\times$ to $1.64\times$, with the strongest advantages on tasks whose structure directly matches the SN’s compositional biases. On Fashion-MNIST (Table 6), SNs achieve competitive classification accuracy, and the stable training of 25-layer networks illustrates that SNs can be optimized at substantial depth when equipped with residual connections and normalization.